Volcanic activity continued at a similar level today although less rockfalls were recorded than has been the case in the last few days. There are indications that rockfall activity has, at least in part, switched to the Tar River Valley side of the dome. The broad band system triggered 35 hybrid earthquakes, 41 LP earthquakes and 39 rock-falls today.
Cloud today prevented good views of the dome but an observation flight reported that ash was being generated to the east of the volcano rather than to the south. The seismic amplitude of rockfalls today was much greater on the Long Ground seismometer relative to the other stations, which also suggests that activity has switched to the eastern side of the dome.
Air quality at all monitoring stations across the island is good today. A GPS survey was carried out today to the West side of the volcano, the results of this survey will be reported at a later date.